OpenStudy (anonymous):

What are the solutions of the quadratic equation? x^2 - 9x + 20 = 0

OpenStudy (owlcoffee):

In this problem, we can observe it's a 2nd degree equality, because the "x" with the highest exponent is the "x2 with a exponent of 2. So, we have to use the general formula: \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] all you have to do is replace the values :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Can you stick around until I get my answer?

OpenStudy (anonymous):

@Owlcoffee Im stuck I got to \[x = -9 \pm \sqrt{1}\]

OpenStudy (owlcoffee):

Oh, you made a little mistake :(

OpenStudy (anonymous):

nooo. help :(

OpenStudy (owlcoffee):

So, any quadratic equation has a general form of: \[ax^2 + bx + c=0\] in order to find the value of x, the general formula comes in action: \[x=\frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\] so, in the problem: \[x^2 -9x +20=0\] as a trick I used back in the day, I separeted them as a sum, so I could see a,b and c more clearly: \[x^2 + (-9)x+20=0\] so, then: \[a=1 , b=-9,c=20\] recplacing in the quadratic formula: \[x=\frac{ -(-9) \pm \sqrt{(-9)^2 -4(1)(20)} }{ 2(1) }\] can you take over from here? :)

OpenStudy (anonymous):

well thats what I did! and I simplified and ended up with \[x = 9 \pm \sqrt{1}\] all over 2

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (anonymous):

and I got stuck there.

OpenStudy (owlcoffee):

good, and what is the square root of 1?

OpenStudy (anonymous):

1

OpenStudy (owlcoffee):

good, so we have: \[x=\frac{ 9 \pm 1 }{ 2 }\] so, in order to get rid of the plus/minus, we have to take both situations: \[x_1 = \frac{ 9 +1 }{ 2 }\] \[x_2 = \frac{ 9-1 }{ 2 }\]